# PSGait: Multimodal Gait Recognition using Parsing Skeleton

### v1

@@ III-C Align Center
Skeleton GaitGraph2(CVPRW2022)[20] 5.0 5.3 5.8 6.2 5.6 11.1 - - -
GaitTR(ES2023)[21] 15.7 18.3 18.5 17.5 17.5 6.6 - - -
Silhouette GaitSet(AAAI2019)[22] 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
GaitPart(CVPR2020)[23] 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
GaitBase(CVPR2023)[18] 71.6 75.0 76.8 78.6 75.5 64.6 - - -
[…]
Skeleton+Silhouette SkeletonGait++†(AAAI2024)[9] 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
[…]
IV Experience
IV-A Datasets
@@
We evaluated our proposed method on three commonly used datasets: SUSTech1K[19] and CCPG[17], which are collected in fully constrained laboratory environments, and Gait3D[8], which is collected from real-world scenarios.
@@
SUSTech1K[19] contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°). The diversity and granularity of the dataset make SUSTech1K[19] an ideal platform for studying the performance of gait recognition in complex environments.
@@
CCPG[17] was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance. The dataset includes 200 identities and more than 16,000 sequences, with each identity captured from 10 viewpoints under seven different clothing change conditions, covering both indoor and outdoor environments. It provides two data modalities: RGB images and silhouette images.
@@
Gait3D[8] is a large-scale gait dataset collected in real-world environments. It was recorded in a supermarket and contains 1,090 hours of video footage with a resolution of 1,920×1,080 and a frame rate of 25 FPS. Gait3D provides four modalities: silhouettes, 2D and 3D skeletons, and 3D meshes.
@@
TABLE III: Impact of circle radius and line width on the SUSTech1K[19] dataset using GaitBase[18].
@@
TABLE IV: Impact of input modalities and fusion strategies on the SUSTech1K[19] dataset using GaitBase[18].
Skeleton Silehouette Overlay Probe Sequence(R-1) Overall
52.5 50.1 35.2 48.4 40.3 51.3 61.1 22.0 50.3 72.5
81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
82.7 81.2 47.8 78.9 81.3 79.8 88.2 35.1 79.7 92.2
86.1 84.5 55.0 82.9 85.4 84.1 90.7 47.1 83.4 94.3
IV-B Implementation Details
For data preprocessing, the original data is used to extract silhouette information from all three datasets. Since skeleton data is not available for the CCPG dataset, we utilize the Sapiens[26] method for pose estimation. The original skeleton data is used for the other two datasets, and the format of points follows COCO17. A confidence threshold of 0.1 is applied.
In the parsing stage, to ensure that the Parsing Skeleton features of the processed image can be effectively captured by the CNN, the radius of the circle representing the head and the width of the line segment representing the limbs are determined based on the image size ratio in each dataset. Specifically, for the SUSTech1K[19] dataset, a radius of 10 and a line width of 12 are used. For the CCPG[17] dataset, the radius and line width are set to 16 and 20, respectively. Similarly, for the Gait3D[8] dataset, the radius and line width are also set to 16 and 20, respectively. The colors selected for parsing each body part are shown in Fig. 2.
To evaluate the versatility of our method, we apply it to two state-of-the-art gait recognition models: GaitBase[18] and DeepGaitV2[25]. The size of the fused image is adjusted to $64\times 44$ before being input into the gait recognition model. For model training, we largely follow the parameter settings from OpenGait[18], an excellent open-source framework in the field of gait recognition. Specifically, for the SUSTech1K[19] dataset, the batch size is set to $(8,8)$, and learning rate milestones are scheduled at 20k, 30k, and 40k steps, with a total of 50k training steps. For the CCPG[17] dataset, we use a batch size of $(8,16)$ with learning rate milestones at 20k, 40k, and 50k steps, and a total of 60k iterations. For the Gait3D[8] dataset, the batch size is set to $(32,4)$, with learning rate milestones occurring at 20k, 40k, and 50k steps, and a total of 60k iterations. The optimization process is configured using the SGD solver. Specifically, the learning rate $lr$ is initialized at 0.1, providing an appropriate step size for updating the model weights during training. A momentum value of 0.9 is employed to accelerate convergence by reducing oscillations and incorporating a weighted history of previous gradients.
IV-C Comparison with State-of-the-art Methods
As shown in Table I and Table II, we evaluate the performance on three datasets. The results highlight the strong competitiveness and versatility of our method.
Competitiveness. We compare and validate the advantages of our method across three datasets. On the SUSTech1K[19] dataset (Table I), compared to SkeletonGait++, Ours+GaitBase improves the Rank-1 accuracy from 81.2% to 83.4% (+2.2%). This demonstrates the effectiveness of our method under the same baseline settings. Furthermore, Ours+DeepGaitV2 achieves the best overall performance on SUSTech1K, attaining a Rank-1 accuracy of 87.9% and a Rank-5 accuracy of 94.8%. On the CCPG[17] dataset (Table II), compared to SkeletonGait++[9], Ours+GaitBase improves the mean accuracy from 82.6% to 86.4% (+3.8%). On the Gait3D[8] dataset (Table II), Rank-1 accuracy is improved from 64.8% to 70.0% (+5.2%). Additionally, Ours+GaitBase and Ours+DeepGaitV2 achieve the best overall performance on CCPG and Gait3D, with a mean accuracy of 86.4% on CCPG and a Rank-1 accuracy of 85.6% and mAP of 69.4% on Gait3D, outperforming all other state-of-the-art methods. These improvements underscore the robustness and adaptability of our approach across diverse benchmarks and evaluation metrics.
Versatility. We further demonstrate the versatility of our approach by integrating it into two baseline models, GaitBase[18] and DeepGaitV2[25], and comparing the resulting performance improvements. On the three datasets mentioned above, our method provides a significant boost over the existing baselines. Specifically, on the SUSTech1K dataset[19], Ours+GaitBase improves the Rank-1 accuracy from 76.1% to 83.4% (+7.3%), while Ours+DeepGaitV2 enhances the Rank-1 accuracy from 77.4% to 84.6% (+7.2%). On the CCPG[17] dataset, integrating our method with GaitBase increases the mean accuracy from 75.5% to 86.4% (+10.9%). For DeepGaitV2, the mean accuracy improves from 83.3% to 85.6% (+2.3%). On Gait3D[8], Ours+GaitBase improves the Rank-1 accuracy from 64.6% to 70.0% (+5.4%). For DeepGaitV2, the Rank-1 accuracy increases from 74.4% to 76.2% (+1.8%), and the mAP rises from 65.8% to 69.4% (+3.6%). These consistent performance improvements across multiple datasets and baseline models highlight the strong versatility of our method. It effectively enhances the performance of any gait recognition model, making it plug-and-play and universally applicable to various scenarios and datasets.
IV-D Ablation Study
@@
To validate the effectiveness of our method, we conducted ablation studies on the SUSTech1K[19] dataset using GaitBase[18] as the baseline gait recognition model. We conducted experiments to assess the impact of key parameters, including circle radius, line width, input modalities, and different data fusion strategies.
Influence of Circle Radius and Line Width. As shown in Table III, we evaluated the impact of circle radius and line width on recognition performance. A smaller setting results in a relatively low Rank-1 accuracy of 82.1%, while a moderate setting achieves the best Rank-1 accuracy of 86.1% and Rank-5 accuracy of 94.3%. However, further increasing the radius and width, slightly reduces Rank-1 accuracy to 84.6%, indicating that finding an optimal balance is crucial for performance. The results indicate that choosing an appropriate circle radius and line width is critical for effectively extracting parsing skeleton features. If these values are too small, the parsing skeleton cannot be effectively captured by the CNN after resizing the image. Conversely, if they are too large, the parsing skeleton may obscure the silhouette, impeding the extraction of human shape features by gait recognition models.
@@
Impact of Data Fusion Strategies. We further analyzed the contribution of different modalities and fusion methods (parsing skeleton, silhouette, and overlay) to overall performance. As shown in Table IV, using parsing skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition. Incorporating silhouette information significantly improves performance, with Rank-1 increasing to 79.7% and Rank-5 reaching 92.2%. In this case, the data fusion strategy employed is a multi-channel fusion method. When parsing skeleton and silhouette are overlaid, the model achieves the best performance, with a Rank-1 accuracy of 83.4% and a Rank-5 accuracy of 94.3%. The overlay fusion strategy capitalizes on the complementary strengths of skeleton and silhouette data, allowing the model to capture both global shape and local structural details. These results demonstrate that our fusion strategy is both effective and achieves an optimal balance between accuracy and robustness.
V Conclusion

### v2

@@ 3.4. Data Augmentation
Skeleton GaitGraph2(Teepe et al., 2022) CVPRW2022 5.0 5.3 5.8 6.2 5.6 11.1 -
GaitTR(Zhang et al., 2023) ES2023 15.7 18.3 18.5 17.5 17.5 6.6 -
Silhouette GaitSet(Chao et al., 2021) AAAI2019 60.2 65.2 65.1 68.5 64.8 36.7 58.3 30.0 17.3
GaitPart(Fan et al., 2020) CVPR2020 64.3 67.8 68.6 71.7 68.1 28.2 47.6 21.6 12.4
GaitBase(Fan et al., 2023b) CVPR2023 71.6 75.0 76.8 78.6 75.5 64.6 -
DeepGaitV2(Fan et al., 2023a) Arxiv2023 78.6 84.8 80.7 89.2 83.3 74.4 88.0 65.8 -
Silhouette+ Skeleton SkeletonGait++(Fan et al., 2024) AAAI2024 77.0 81.4 83.0 88.8 82.6 64.8 80.5 54.8 36.1
PSGait-B Ours 82.5 85.3 88.6 89.1 86.4 70.0 85.7 62.2 43.3
PSGait-D Ours 80.9 84.8 86.4 90.1 85.6 76.2 88.7 69.4 50.8
Table 2. Quantitative comparison of gait recognition methods on the SUSTech1K(Shen et al., 2023).
Input Method Publication Probe Sequence (R-1) Overall
Skeleton GaitGraph2(Teepe et al., 2022) CVPRW2022 22.2 18.2 6.8 18.6 13.4 19.2 27.3 16.4 18.6 40.2
GaitTR(Zhang et al., 2023) ES2023 33.3 31.5 21.0 30.4 22.7 34.6 44.9 23.5 30.8 56.0
Silhouette GaitSet(Chao et al., 2021) AAAI2019 69.1 68.2 37.4 65.0 63.1 61.0 67.2 23.0 65.0 84.8
GaitPart(Fan et al., 2020) CVPR2019 62.2 62.8 33.1 59.5 57.2 54.8 57.2 21.7 59.2 80.8
GaitGL(Lin et al., 2022) ICCV2021 67.1 66.2 35.9 63.3 61.6 58.1 66.6 17.9 63.1 82.8
GaitBase(Fan et al., 2023b) CVPR2023 81.5 77.5 49.6 75.8 75.5 76.7 81.4 25.9 76.1 89.4
DeepGaitV2(Fan et al., 2023a) Arxiv2023 83.5 79.5 46.3 76.8 79.1 78.5 81.1 27.3 77.4 90.2
Silhouette+ Skeleton BiFusion(Peng et al., 2024) MTA2023 69.8 62.3 45.4 60.9 54.3 63.5 77.8 33.7 62.1 83.4
SkeletonGait++(Fan et al., 2024) AAAI2024 84.7 82.4 46.7 81.8 81.2 82.6 85.1 44.3 81.2 93.2
PSGait-B Ours 86.2 84.5 55.1 83.0 85.4 84.2 90.7 47.2 83.4 94.3
PSGait-D Ours 87.9 85.2 53.9 84.5 86.3 85.3 91.4 47.5 84.6 94.8
3.5. Gait Recognition Model
The current gait recognition framework employs a backbone network $\mathcal{B}$ to extract features $\mathbf{f}$ from an input sequence of gait images $\mathbf{s}$, defined as:
$\mathbf{f}=\mathcal{B}(\mathbf{s})\in\mathbb{R}^{n\times c\times s\times h\times w}$
where $n$, $c$, $s$, $h$, and $w$ denote the number of gait frames, the number of channels, the sequence length, and the height and width of each image, respectively. Temporal Pooling (TP) (Chao et al., 2021) is applied to compress these features into more compact representations $z$:
$z=TP({\mathbf{f}})\in\mathbb{R}^{n\times c\times h\times w}$
Subsequently, Horizontal Pooling (HP) (Fu et al., 2019) divides these feature maps into horizontal segments, pooling each segment into vectors $z_{s}$ to enable efficient feature extraction through global pooling across these segments:
$\mathbf{f}^{\prime}=\text{maxpool}(z_{s})+\text{avgpool}(z_{s})$
A fully connected layer then projects these pooled vectors into a metric space, followed by Batch Normalization Neck (BNNeck) (Luo et al., 2019) to refine the feature space. The final classification is performed via cross-entropy loss:
$L_{ce}=-\sum_{i=1}^{n}y_{i}\log(\widehat{y_{i}})$
where $y_{i}$ denotes the true label and $\widehat{y_{i}}$ represents the predicted probability for each class.
Additionally, the model employs triplet loss to enhance the discrimination of subtle gait variations essential for accurate classification. During training, $N$ triplets are formed in each batch, where each triplet consists of an anchor sequence $\mathbf{s}_{i}^{a}$, a positive sequence $\mathbf{s}_{i}^{p}$ from the same identity, and a negative sequence $\mathbf{s}_{i}^{n}$ from a different identity. The triplet loss(Schroff et al., 2015) is computed as follows:
$L_{triplet}=\sum_{i=1}^{N}\max\left(0,\|f(\mathbf{s}_{i}^{a})-f(\mathbf{s}_{i}^{p})\|_{2}^{2}-\|f(\mathbf{s}_{i}^{a})-f(\mathbf{s}_{i}^{n})\|_{2}^{2}+\alpha\right)$
where $f(\cdot)$ is the embedding function that maps each sequence to the embedding space, $|\cdot|_{2}$ represents the Euclidean distance, and $\alpha$ is a margin parameter that enforces a separation between positive and negative pairs. This strategy improves the model’s discriminative ability between classes.
The overall objective function is a combination of the cross-entropy loss and the triplet loss, formulated as:
$L_{total}=L_{ce}+L_{triplet}$
This integrated loss function optimizes the gait recognition model, enhancing its accuracy in distinguishing between different identities based on gait features.
Table 3. Generalization experiments on two baseline models across three datasets.
Method SUSTech1K Gait3D CCPG
rank-1 rank-5 rank-1 mAP rank-1
GaitBase(Fan et al., 2023b) 76.1 89.4 64.6 54.5 75.5
GaitBase+Parsing Skeleton 83.4(7.3 $\uparrow$) 94.3(4.9 $\uparrow$) 70.0(5.4 $\uparrow$) 62.2(7.7 $\uparrow$) 86.4(10.9$\uparrow$)
DeepGaitV2(Fan et al., 2023a) 77.4 90.2 74.4 65.8 83.3
DeepGaitV2+Parsing Skeleton 84.6(7.2 $\uparrow$) 94.8(4.6 $\uparrow$) 76.2(1.8 $\uparrow$) 69.4(3.6$\uparrow$) 85.6(2.3$\uparrow$)
Table 4. Efficiency and accuracy comparison between PSGait-B and the SOTA method.
Method Parameters GPU memory Preprocessing time Training time SUSTech1K CCPG
(M) (GB) (ms per frame) (s per epoch) rank-1 rank-1
GaitBase(Fan et al., 2023b) 8.02 4.68 0 0.22 76.1 75.5
SkeletonGait++(Fan et al., 2024) 9.13 5.14 0.71 0.34 81.3 83.7
PSGait-B(Ours) 8.02(12$\downarrow$) 4.70(9$\downarrow$) 0.34(52$\downarrow$) 0.27(21$\downarrow$) 83.4(2$\uparrow$) 86.4(3$\uparrow$)
4. EXPERIMENTS
4.1. Datasets
@@
We evaluated our proposed method on three commonly used datasets: SUSTech1K(Shen et al., 2023) and CCPG(Li et al., 2023b), which are collected in fully constrained laboratory environments, and Gait3D(Zheng et al., 2022), which is collected from real-world scenarios. Our experiments strictly follow the official evaluation protocols.
@@
SUSTech1K(Shen et al., 2023) contains gait data from 1,050 subjects, covering 8 different conditions: normal (NM), bag (BG), clothes changing (CL), carrying (CR), umbrella (UB), uniform (UF), occlusion (OC), and night (NT), as well as 12 viewpoints (ranging from 0° to 315°). The diversity and granularity of the dataset make SUSTech1K(Shen et al., 2023) an ideal platform for studying the performance of gait recognition in complex environments.
@@
CCPG(Li et al., 2023b) was designed to address the challenge of gait recognition under clothing change conditions and is particularly suitable for dynamic person recognition in video surveillance. The dataset includes 200 identities and more than 16,000 sequences, with each identity captured from 10 viewpoints under seven different clothing change conditions, covering both indoor and outdoor environments. It provides two data modalities: RGB images and silhouette images.
@@
Gait3D(Zheng et al., 2022) is a large-scale gait dataset collected in real-world environments. It was recorded in a supermarket and contains 1,090 hours of video footage with a resolution of 1,920×1,080 and a frame rate of 25 FPS. Gait3D provides four modalities: silhouettes, 2D and 3D skeletons, and 3D meshes.
4.2. Implementation Details
For data preprocessing, silhouettes are extracted from all three datasets. Since skeleton data is unavailable for CCPG, we employ Sapiens (Khirodkar et al., 2024) for 2D pose estimation, while original skeleton annotations are used for SUSTech1K and Gait3D in COCO17(Lin et al., 2014) format, with a confidence threshold of 0.1.
In the parsing stage, to ensure the Parsing Skeleton is clearly captured by CNNs, the circle radius (for the head) and line width (for limbs) are adjusted according to image size. Specifically, SUSTech1K uses radius 10 and width 12; CCPG and Gait3D use radius 16 and width 20. The color scheme for each body part is shown in Fig. 2.
To evaluate generalization, Parsing Skeletons are integrated into two state-of-the-art gait recognition models: GaitBase(Fan et al., 2023b) and DeepGaitV2(Fan et al., 2023a). The fused inputs are resized to $64\times 44$ before being fed into the recognition models.
For training, we almost follow the OpenGait (Fan et al., 2023b) framework. On SUSTech1K(Shen et al., 2023), batch size is set to (8, 8), with learning rate milestones at 20k, 30k, and 40k steps, and a total of 50k steps. On CCPG(Li et al., 2023b), we use (8, 16) with milestones at 20k, 40k, and 50k (60k total). On Gait3D(Zheng et al., 2022), the batch size is (32, 4), also trained for 60k steps. We use SGD with an initial learning rate of 0.1 and momentum of 0.9 to facilitate stable convergence.
@@ 4.2. Implementation Details
Table 5. Impact of circle radius and line width on the SUSTech1K(Shen et al., 2023) dataset using GaitBase(Fan et al., 2023b).
@@
Table 6. Impact of input modalities and fusion strategies on the CCPG(Li et al., 2023b) dataset using GaitBase(Fan et al., 2023b).
Skeleton Silehouette Multi-Channel Overlay-Based CCPG
CL UP DN BG Mean
53.8 58.5 61.5 56.0 57.4
76.5 80.5 82.3 86.4 81.4
79.3 83.3 85.2 86.5 83.6
82.5 85.3 88.6 89.1 86.4
4.3. Comparison with State-of-the-art Methods
We evaluate the proposed PSGait framework against state-of-the-art gait recognition methods across three input modalities-skeletons, silhouettes, and their fusion—on three benchmarks: CCPG(Li et al., 2023b), SUSTech1K(Shen et al., 2023), and Gait3D(Zheng et al., 2022). PSGait-B and PSGait-D adopt GaitBase (Fan et al., 2023b) and DeepGaitV2 (Fan et al., 2023a) as gait recognition models, respectively. In order to highlight the advantages of the Parsing Skeleton representation itself, rather than the influence of the network architecture, we adjusted the downstream settings of SkeletonGait++ to align with those of GaitBase.
On CCPG and Gait3D (Table 1), PSGait consistently outperforms existing methods in both unimodal and multimodal settings. Under the skeleton-only modality, methods like GaitGraph2 (Teepe et al., 2022) and GaitTR (Zhang et al., 2023) achieve low recognition rates due to the sparsity of keypoints. SkeletonGait++ (Fan et al., 2024) improves performance by using Skeleton Maps, but still struggles in real-world scenarios. In contrast, when our Parsing Skeleton is fused with silhouettes, PSGait-B and PSGait-D set new state-of-the-art results, achieving up to 86.4% Rank-1 on CCPG and 76.2% on Gait3D. PSGait-D also brings large gains in mAP and mINP (+12.7% and +14.7%) over the strongest baseline, indicating improved retrieval robustness.
On SUSTech1K (Table 2), PSGait again demonstrates superior performance under various challenging conditions. When our Parsing Skeleton is fused with silhouettes, PSGait-B and PSGait-D consistently outperform SkeletonGait$+$ across all probe types. For example, PSGait-D achieves 91.4% Rank-1 under occlusion and 86.1% overall Rank-1 accuracy, surpassing the baseline by large margins—+7.6% in night and +7.1% in carrying conditions.
These results validate that: (1) Parsing Skeleton provides a richer and more discriminative representation than traditional silhouettes or skeletons; (2) PSGait effectively captures both part-level dynamics and global structure, enabling consistent SOTA performance across datasets; (3) PSGait is accurate, lightweight, and robust, making it highly suitable for real-world gait recognition applications.
4.4. Generalization of the Parsing Skeleton
To evaluate the generalization capability of the proposed Parsing Skeleton, we integrate it into two representative baselines—GaitBase and DeepGaitV2—and test on three diverse datasets: SUSTech1K, Gait3D, and CCPG (Table 3).
On SUSTech1K, Parsing Skeleton brings notable Rank-1 improvements of +7.3% for GaitBase and +7.2% for DeepGaitV2, with consistent gains in Rank-5, demonstrating enhanced retrieval performance. On the challenging Gait3D dataset, GaitBase+Parsing Skeleton improves by +5.4% in Rank-1 and +7.7% in mAP, while DeepGaitV2 gains +1.8% and +3.6%, respectively—showing its effectiveness under real-world, multi-view conditions. On CCPG, Parsing Skeleton improves GaitBase by a substantial +10.9% and DeepGaitV2 by +2.3% in Rank-1, highlighting its robustness across varied environments.
These results confirm that Parsing Skeleton is a model-agnostic, plug-and-play, generalizable representation. It consistently improves performance across architectures and datasets, making it well-suited for real-world deployment in gait recognition systems.
4.5. Lightweight of the PSGait
To evaluate computational efficiency, we compare PSGait-B with two baselines—GaitBase(Fan et al., 2023b) (silhouette-only) and SkeletonGait++(Fan et al., 2024) (skeleton map + silhouette)—in terms of model size, memory usage, preprocessing/training time, and accuracy. All evaluations were conducted on 100,000 frames from SUSTech1K(Shen et al., 2023), using GaitBase as the backbone and four RTX 2080 GPUs.
As shown in Table 4, PSGait-B achieves a superior trade-off between efficiency and performance. It reduces parameters by 12%, GPU memory by 9%, preprocessing time by 52%, and training time by 21% compared to SkeletonGait++. Despite these reductions, PSGait-B achieves +2% Rank-1 accuracy on SUSTech1K and +3% on CCPG(Li et al., 2023b), showing better performance at lower cost.
Compared to GaitBase, PSGait-B maintains similar resource usage but brings substantial accuracy gains (+7.3% on SUSTech1K, +10.9% on CCPG), validating that the Parsing Skeleton effectively enhances model capacity without increasing complexity.
In summary, PSGait is a lightweight and efficient solution that delivers high performance and is well-suited for real-time applications or limited-resource environments.
4.6. Ablation Study
@@
To validate the effectiveness of our method, we conducted ablation studies on SUSTech1K(Shen et al., 2023) and CCPG(Li et al., 2023b) using GaitBase(Fan et al., 2023b) as the baseline gait recognition model. We conducted experiments to assess the impact of key parameters, including circle radius, line width, input modalities, and different data fusion strategies.
Circle Radius and Line Width. As shown in Table 5, we evaluated the impact of circle radius and line width on recognition performance. A smaller setting leads to relatively lower accuracy, while a moderate setting yields the best performance in terms of Rank-1 accuracy. However, increasing the radius and width beyond this point results in a slight decrease in Rank-1 accuracy. Smaller values lead to under-represented structures, while larger ones may obscure silhouette contours, degrading performance. This highlights the importance of carefully balancing visual clarity and structural integrity when rendering Parsing Skeletons.
@@
Fusion Strategies. Table 6 compares different input modalities and fusion approaches. Here, “Skeleton” refers to Parsing Skeleton. Using Parsing Skeletons or silhouettes alone does not yield particularly high performance, indicating that relying on either type of data alone is insufficient for accurate gait recognition. Incorporating skeleton and silhouette information improves performance, with Rank-1 increasing to 83.6%. The best performance is achieved by overlaying Parsing Skeletons onto silhouettes. This overlay strategy effectively integrates part-level dynamics with global shape cues, enabling more robust and discriminative gait representations.
These results confirm that both visual rendering parameters and fusion strategy are crucial to fully unlocking the potential of Parsing Skeleton for gait recognition.
5. CONCLUSION